OpenStudy (anonymous):

Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions. x2 = 5x + 6 Two different imaginary solutions Two different rational solutions Two different irrational solutions Exactly one rational solution

OpenStudy (anonymous):

The discriminant is: b^2-4ac Now, what are: a=? b=? c=? Do you know how to do this ?

OpenStudy (anonymous):

you first need to write the equation in this form: ax^2+bx+c=0

OpenStudy (chaise):

a = 1 b = -5 c = -6 b^2-4ac = 25-4x1x-6 = 49 The equation has two REAL anwers.

OpenStudy (anonymous):

Two different imaginary solutions Two different rational solutions Two different irrational solutions Exactly one rational solution which of of these are correct?

OpenStudy (anonymous):

The discriminant = 49 which is a perfect square, so: Two different rational solutions

OpenStudy (anonymous):